# Broken Expectations: The Effects of Modelling Assumptions on the Inferred Dark Matter Distribution in the Milky Way’s Satellites

### v1

@@ 2.4 J-Factor Calculation
with the distance along the line-of-sight, which runs from 0 to , and is the solid angle over which we integrate. In our study, we fix this to 0.5∘ as done in most J-factor studies and corresponds for most dwarfs to a radius likely to contain most of the annihilation signal (Calore et al., 2018).
In the WIMP scenario, it is expected that dark matter annihilation can lead to the emission of gamma rays. Thus the J-factor enters the differential gamma-ray flux through the following equation:
where is the mass of the dark matter particle, is the differential number of photons per annihilation, and is the velocity-averaged annihilation cross-section (Cirelli et al., 2011). For a given gamma-ray production channel, if the J-factor is known and differential gamma-ray flux is measured, constraints can be placed on the mass and annihilation cross-section of dark matter.
To re-parameterize the density function in terms of and (or in the case of spherical symmetry only ), we use the method of Geringer-Sameth et al. (2015), taking a new variable which denotes the distance along the line of sight, from the point where the line of sight makes its closest approach to the centre of the spherical system. This means that the line of sight with an angular separation of has an impact parameter , with being the distance of the spherical system’s centre to the observer. Thus becomes and limits of integration for become to . Furthermore one can simply replace .
In this work, we calculate J-factors in two ways. The first is through evaluating Equation 19 with the density profiles recovered by pyGravSphere (i.e. observationally-derived J-factors). The second method computes the “true" J-factor, where the density of the simulated dSph is estimated by calculating the local density of each bound dark matter particle. For each particle we find nearest neighbours and calculate the density inside of a sphere containing all neighbours, centred on the particle . For the density estimation, we select only the particles within the 0.5∘ cone starting from the observer and extending to a distance D to the dwarf. To optimize the search for the nearest neighbours, we use a KD-tree method (Bentley, 1975). The densities are then summed inside the solid angle, using the following equation to approximately obtain the final “true" J-factor (Kuhlen, 2010):
@@ 3.1 Wolf Mass Estimator
On the left of Figure 5, we summarise the results of the application of the 2-D Wolf mass estimator on our sample of dwarfs in the 2 Milky Way potentials, obtained through fitting surface brightness profiles to our dwarfs. We display the results for the present day as well as the nearest peri and apocentres. It is clear that in the ‘light’ potential, the mass within the half-mass radius is typically overestimated, while in the ‘heavy’ potential the mass is typically underestimated by 5-10 per cent. Exceptions to this include Ursa Minor, where both in the ‘heavy’ and the ‘light’ potential, the mass is underestimated. For Draco, Sculptor and Ursa Minor – the dwarfs which come closest to the Milky Way – we see some discrepancies in the accuracy of masses estimated at peri and apocentre. For mass estimates at the present day, the true value lies within the 68 per cent confidence levels in both potentials for all dwarfs but Ursa Minor in the heavy potential.
@@
To understand the origin of the discrepancy in the results obtained at peri- and apocentres, we examine the extreme case of Ursa Minor in a ‘heavy’ Milky Way potential over the whole course of its orbital history. This is shown in Figure 6 , where results for both dark matter and stellar tracers, as well as the 3-D and 2-D estimators, are shown.
3.1.1 Dark matter in Ursa Minor
@@
It is worth noting that the velocity dispersion profiles of NFW haloes are expected to rise, breaking the assumption of nearly flat velocity dispersion of the Wolf et al. (2010) mass estimator. Nevertheless, we proceed with our analysis under the assumption that it causes a constant bias in the performance of the estimator on a dark matter halo. We immediately see an interesting oscillating effect for the two 2-D estimators and one less pronounced for the 3-D estimator in the top panel of Figure 6, corresponding to Ursa Minor’s orbit from pericenter to apocenter and back. In the dark matter, this effect is clear for all of the dSphs. This shows that tidal effects play an important role in the accuracy of the mass estimates of these galaxies. We note that the oscillations seem to follow a clear pattern with the dSph orbit, with the pattern in the 2-D mass ratio computed with surface density profile fitting appearing more complex with additional peaks and troughs. We also note that the first few Gyr show a large inaccuracy in the estimator, which we attribute to a numerical artefact of the simulation since we suddenly plunge our dSphs into an orbit around the Milky Way potential.
@@
To understand the origin of these oscillating patterns, at the bottom of Figure 6 we show the ratio of the minor to the major axis of the dark matter halo as a function of time. We compute the principal axes using the eigenvalues of the inertia tensor (Bett et al., 2007). The blue line shows the ratio of the 3-D major and minor axes, while the orange line shows the ratio of the 2D on-the-sky minor and major axes. It is evident that spherical symmetry is broken both in the 3-D and 2-D cases, again in an oscillatory fashion corresponding to the orbit of Ursa Minor, with the dark matter halo shape becoming more stretched out at pericentre, leading the ‘peaks’ in the accuracy of the 3-D mass estimator. Over time, the amplitude of the stretching becomes smaller as the dwarf becomes more spherical and the amplitude of the oscillations in the mass estimator accuracy also appears to begin converging.
@@
The 2-D axis ratio also demonstrates overall stretching when Ursa Minor is at pericentre, but not always. At every second pericentre, the 2-D axis ratio seems to change quickly from stretched-out to nearly perfect sphericity (when the dwarf is clearly deformed in 3-D). This indicates that in some instances, the 3-D major axis of the dwarf is aligned closely with the line-of-sight, thus causing the dSph to appear spherical. From the top panel, it can be seen that both the stretching in the plane of the sky and the times when the dwarf appears artificially spherical cause peaks in the estimator’s accuracy. Perhaps surprisingly, the mass estimated at pericentre can be more accurate, though this effect is artificial as, on average, the 2-D estimator is biased low by 15 per cent for the dark matter component in Ursa Minor. Over time, as the dwarf becomes more spherical, these peaks decrease in amplitude, while the estimator’s accuracy slowly increases with increased accuracy at apocentre. This is a consistent pattern we find for the dwarfs in our sample.
@@
Finally, we compare the result again with the modified 2-D Wolf estimator, which uses a directly obtained half-mass radius and mean line-of-sight velocity dispersion within this radius, shown in red in Figure 6. In this case, we see that the estimator accuracy oscillates resonantly with the orbital phase, with more accurate results at apocentre and less accurate at pericentre.
These results raise an important question of whether it is the change in the dwarf’s shape due to tides or some fundamental lack of equilibrium that is causing the pattern in the estimator’s accuracy with peri and apocentres. Evidently, both effects are at play as the 2-D estimator returns higher masses when the dwarf appears more spherical and the pattern in the 3-D estimator persists even when the dwarfs become more spherical. Moreover, we find that the amplitude of the oscillations is related to the orbital eccentricity, with more prominent peaks for more eccentric orbits, while the asphericity induced by tides has a similar magnitude for all of our dwarfs. This is perhaps clearest from the results of the 2-D estimator when using the velocity dispersion within the half-light radius (red line), where the pattern does not follow as clearly as that of the dwarf’s on-the-sky ellipticity, with only minor ‘peaks’ in estimated mass at pericentre, where the dwarf appears more spherical.
3.1.2 Stars in Ursa Minor
@@
In the middle panel of Figure 6, we show similar results for the stars in Ursa Minor. The 3-D Wolf estimate is accurate to 5-10 per cent and displays a weak trend with the dwarf spheroidal orbit. It should be noted that the mass estimates are very noisy due to low numbers of stellar particles, which are nevertheless comparable to the largest kinematic samples available for the Milky Way’s dSphs (Read et al., 2019). The oscillating pattern related to the orbital phase can be seen more clearly in the 2-D mass estimator. We should note that such a pattern is only obvious in our sample for dwarfs on particularly eccentric orbits, with eccentricity 0.5: Ursa Minor in the heavy potential, Draco in the light and heavy potential and Sculptor in the light potential, which is likely a result of more severe tidal shocks in these objects (Gnedin et al., 1999). In the lower panel of Figure 6, only minor changes in the projected axes are present888We note that for none of our simulated dwarfs does the 2-D axis ratio decrease beyond 0.1 (ellipticity ), making it unlikely that “stretching” due to tides can alone explain the observed ellipticities of these dwarf spheroidal systems (Iorio et al., 2019)., and the 2-D mass estimator is accurate to per cent. This is not surprising, as the stellar component is significantly more gravitationally bound than the dark matter, has a shorter dynamical time, and is, therefore, less likely to be deformed by the action of tides. It is clear from Figure 6, however, that the mean accuracy of the Wolf estimator on the stellar tracers is affected by the mass of the Milky Way. We discuss this phenomenon in the following.
@@
To understand what causes the systematic underestimation of the mass within the half-mass radius in the ‘heavy’ potential, in Figure 7 we display the line-of-sight velocity dispersion profiles computed for our sample of simulated dwarfs in the two potentials. Immediately, it is clear that in the ‘light’ Milky Way potential, the line-of-sight velocity dispersion profile typically rises towards the outskirts of the galaxy (which is consistent with the kinematics of a Plummer distribution in an NFW halo in our initial conditions999We note that several sources in the literature report that rising line-of-sight velocity dispersion can be interpreted as the consequence of tides (Johnston et al., 1999; Read et al., 2006; Muñoz et al., 2008) due to inclusion of unbound stars in the tidal tails, however in our initial conditions this is simply a consequence of modelling a Plummer stellar distribution in an NFW halo (see also Stoehr et al. 2002) and is particularly prominent in Fornax, Carina and Sculptor (see the Appendix). As a consequence of tides, the line-of-sight velocity dispersion profiles in our simulations tend to flatten and, in a heavy ‘potential’, drop towards the outskirts of the galaxy.), while in the ‘heavy’ Milky Way potential, the line-of-sight velocity dispersion profile tends to drop in the outskirts. Therefore, when computing the luminosity-weighted line-of-sight velocity dispersion through Equation 16 in the ‘light’ potential, this tends to produce a larger velocity dispersion, while the assumption of a nearly constant velocity dispersion profile in the Wolf et al. (2010) estimator is broken. In contrast, in the ‘heavy’ potential, the line-of-sight velocity dispersion estimate is typically lower. The use of the simplified form of the 2-D mass estimator, with velocity dispersion below the half-light radius, results in selecting parts of the velocity dispersion profile which are typically flat (aside from Draco), thus leading to lower velocity dispersion measurements and masses in the ‘light’ potential and higher in the ‘heavy’ potential. This trend is particularly prominent at dwarf apocentres in Figure 5, where the dwarfs are typically more spherical in 2-D and 3-D, thus minimising the effects of any inaccuracies when estimating the half-light radius, or loss of equilibrium at pericentre.
@@
This opens an interesting avenue to explore constraining the mass of the Milky Way by jointly analysing the line-of-sight velocity dispersion profiles of its satellite dwarf galaxies. However, it would be of essence to obtain data beyond the half-light radii of the dwarf satellites, into their stellar halos (Jensen et al., 2024). We caution, however, that these conclusions also depend on the internal shape of the galaxies, their stellar light distribution (for a fixed model of dark matter), and full tidal history (since orbits of many dwarfs are affected by the presence of the LMC).
@@ 3.1.4 Discussion
Although we have simulated spherical systems, we found that tidal forces can cause deformations in the shape of the simulated dwarfs. This effect, which is the most significant at pericentre, is obvious in the dark matter but can also be seen in the stars, especially for eccentric dwarf galaxy orbits. While the estimator recovers true within its errors in most cases and is typically accurate within 10 per cent, our findings suggest the results of the estimator should be interpreted with care and to bear in mind the orbital properties and the current orbital phase of the subject galaxy. Additionally, it is unclear how the estimator would behave if the subject galaxy is aspherical and acted upon by tides. Our experiment with using dark matter as tracers demonstrates that for a population of tracers that extends to the tidal radius, the performance of the estimator tends to degrade, with mass typically underestimated by 15-20 per cent until the halo stops losing significant fractions of mass, becomes more spherical, and settles towards the final bound remnant state (Errani & Navarro, 2021).
@@ 3.2 Results with pyGravSphere
Fig. 8 show the results for Carina and Fornax, which have relatively low orbital eccentricity in both potentials, with orbits close to circular. For this reason, these galaxies have bound stars extending further out than the rest of our sample, with 8 kpc for Fornax and 2 kpc for Carina. While this is further out than the samples of data used by Read et al. (2019), in Section C of the Appendix, we will show that limiting the maximum radius of data produces consistent results. For Fornax in the ‘light’ and ‘heavy’ potential, the density is underestimated in the inner regions and overestimated near the half-light radius (black dashed line). This introduces a pattern that we will see in the remainder of our results. We note that such behaviour has been seen in the GravSphere method before when modelling Plummer-like stellar distributions inside cuspy haloes, especially with radially rising anisotropy (Read & Steger, 2017; Read et al., 2021). The origin of this discrepancy is not fully clear, but we have confirmed that this is not due to tidal effects modifying the outer profile. We have tested this by modifying the priors of the broken power-law model to allow slopes as steep as . While this improved the fit to the outer profile in dwarfs that are particularly affected by tides, the fit to the inner profile did not improve and the bias near the half-light radius remained. Truncating the sample of stellar particles to the inner 2 kpc in the case of Fornax worsened the fit in the ‘light’ potential, with the bias in the inner slope remaining. In the ‘heavy’ potential, however, this improved the recovery of the central density. We also found that this effect is unlikely due to the disequilibrium of the dwarfs, as this result persists whether the dwarf is modelled at pericentre, apocentre, or in isolation. We discuss the implications of this result in Section 3.3. In Appendix C and D, we discuss the origin of the bias, which is related to the constraint on the “smoothness” of the power-law density profile combined with the use of VSPs.

### v2

@@ 2.4 J-Factor Calculation
with the distance along the line-of-sight, which runs from 0 to , and is the solid angle over which we integrate. In our study, we fix this to 0.5 degrees as done in most J-factor studies and corresponds for most dwarfs to a radius likely to contain most of the annihilation signal (Calore et al., 2018).
In this work, we calculate J-factors in two ways. The first is through evaluating Equation 19 with the density profiles101010To re-parameterize the density distribution, , in terms of and (or in the case of spherical symmetry only ), we use the method of Geringer-Sameth et al. (2015). recovered by pyGravSphere (i.e. observationally-derived J-factors). The second method computes the “true" J-factor, where the density of the simulated dSph is estimated by calculating the local density of each bound dark matter particle. For each particle we find nearest neighbours and calculate the density inside of a sphere containing all neighbours, centred on the particle . For the density estimation, we select only the particles within the 0.5∘ cone starting from the observer and extending to a distance D to the dwarf. To optimize the search for the nearest neighbours, we use a KD-tree method (Bentley, 1975). The densities are then summed inside the solid angle, using the following equation to approximately obtain the final “true" J-factor (Kuhlen, 2010):
@@ 3.1 Wolf Mass Estimator
On the left of Figure 5, we summarise the results of the application of the 2-D Wolf mass estimator on our sample of dwarfs in the 2 Milky Way potentials, obtained through fitting surface brightness profiles to our dwarfs. We display the results for the present day as well as the nearest peri- and apocentres. It is clear that in the ‘light’ potential, the mass within the half-mass radius is typically overestimated, while in the ‘heavy’ potential the mass is typically underestimated by 5-10 per cent. Exceptions to this include Ursa Minor, where both in the ‘heavy’ and the ‘light’ potential, the mass is underestimated. For mass estimates at the present day, the true value lies within the 68 per cent confidence levels in both potentials for all dwarfs but Ursa Minor in the heavy potential.
@@
While the results of the estimator are consistent within their errors for all orbital phases of the dwarfs, it can be seen that the overall accuracy of the estimator can change with orbital phase. To understand this result, we examine the extreme case of Ursa Minor in a ‘heavy’ Milky Way potential over the whole course of its orbital history, applying the Wolf estimator to both dark matter and stars used as tracers.
3.1.1 Dark matter particles as tracers in Ursa Minor
@@
We begin by considering the performance of the estimator when applied to the more spatially extended population of dark matter particles as tracers. The velocity dispersion profiles of NFW haloes are expected to rise due to typically isotropic orbits in the centre found in CDM simulations (Navarro et al., 2010), breaking the assumption of nearly flat velocity dispersion of the Wolf et al. (2010) mass estimator. We thus proceed under the assumption that this causes a constant bias in the performance of the estimator. In the top panel of Figure 6, we see an interesting oscillating effect for the two 2-D estimators and one less pronounced for the 3-D estimator, corresponding to Ursa Minor’s orbit. The oscillations seem to follow a clear pattern with the orbit, with the pattern in the 2-D mass ratio computed with surface density profile fitting (orange) appearing more complex. We also note that the first few Gyr show a large inaccuracy in the estimator, which we attribute to a numerical artefact of the simulation since we suddenly plunge our dSphs into an orbit around the Milky Way potential.
@@
To understand the origin of these oscillations, at the bottom of Figure 6 we show the ratio of the minor to the major axis of the dark matter halo as a function of time. We compute the principal axes using the eigenvalues of the reduced inertia tensor (Bett et al., 2007). The blue line shows the ratio of the 3-D major and minor axes, while the orange line shows the ratio of the 2-D on-the-sky minor and major axes. It is evident that spherical symmetry is broken both in the 3-D and 2-D cases with the orbit of Ursa Minor. The dark matter halo becomes more stretched out at pericentre, leading to ‘peaks’ in the accuracy of the 3-D mass estimator. Over time, the amplitude of the stretching becomes smaller as the halo becomes more spherical and the amplitude of the oscillations in the mass estimator accuracy appears to begin converging.
@@
The 2-D axis ratio also demonstrates overall stretching of the dark matter halo when Ursa Minor is at pericentre, but not always. At every second pericentre, the 2-D axis ratio changes quickly from stretched-out to nearly perfect sphericity (when the dwarf is clearly deformed in 3-D). This indicates that in some instances, the 3-D major axis of the halo is aligned closely with the line-of-sight, thus causing the halo to appear spherical in 2-D. From the top panel, it can be seen that both the stretching in the plane of the sky and the times when the halo appears artificially spherical cause peaks in the estimator’s accuracy.
@@
Finally, we compare the result again with the modified 2-D Wolf estimator, which uses a directly obtained half-mass radius and mean line-of-sight velocity dispersion within this radius, shown in red in Figure 6. In this case, we see that the estimator accuracy oscillates resonantly with the orbital phase, with more accurate results at apocentre and less accurate at pericentre. Only minor fluctutations in the accuracy of the modified 2-D estimator are observed at pericentre, indicating that the corresponing peaks in the 2-D and 3-D estimators stem primarily from effects of tides on the outer parts of the halo, suggesting that these effects are less likely to affect the more embedded stellar population, as we discuss in the following.
3.1.2 Wolf estimator applied to stars in Ursa Minor
@@
In the middle panel of Figure 6, we show the results for applying the estimator to stellar tracers in Ursa Minor. The 3-D Wolf estimate is accurate to 5-10 per cent and displays a weak trend with the dwarf spheroidal orbit. It should be noted that the mass estimates are very noisy due to low numbers of stellar particles, which are nevertheless comparable to the largest kinematic samples available for the Milky Way’s dSphs (Read et al., 2019). The oscillating pattern related to the orbital phase can be seen more clearly in the 2-D mass estimator. The amplitude of the oscillations, even in the extreme case of Ursa Minor, is very small and a pattern is only present for dwarfs on particularly eccentric orbits, with eccentricity 0.5: Ursa Minor in the heavy potential, Draco in the light and heavy potential and Sculptor in the light potential, which is likely a result of more severe tidal shocks in these objects (Gnedin et al., 1999). In the lower panel of Figure 6, only minor changes in the projected axes are present111111We note that for none of our simulated dwarfs does the 2-D axis ratio decrease beyond 0.1 (ellipticity ), making it unlikely that “stretching” due to tides can alone explain the observed ellipticities of these dwarf spheroidal systems (Iorio et al., 2019)., and the 2-D mass estimator is accurate to per cent. This is not surprising, as the stellar component is significantly more gravitationally bound than the dark matter, has a shorter dynamical time, and is, therefore, less likely to be deformed by the action of tides. It is clear from Figure 6, however, that the mean accuracy of the Wolf estimator on the stellar tracers is affected by the mass of the Milky Way. We discuss this phenomenon in the following.
@@
To understand what causes the systematic underestimation of mass in the ‘heavy’ potential, in Figure 7 we display the line-of-sight velocity dispersion profiles computed for our sample of simulated dwarfs in the two potentials. It is clear that in the ‘light’ Milky Way potential, the line-of-sight velocity dispersion profile typically rises towards the outskirts of the galaxy (which is consistent with the kinematics of a Plummer distribution in an NFW halo in our initial conditions), while in the ‘heavy’ Milky Way potential, the line-of-sight velocity dispersion profile tends to drop in the outskirts. Therefore, when computing the luminosity-weighted line-of-sight velocity dispersion through Equation 16 in the ‘light’ potential, this tends to produce a larger velocity dispersion, while the assumption of a nearly constant velocity dispersion profile in the Wolf et al. (2010) estimator is broken. In contrast, in the ‘heavy’ potential, the line-of-sight velocity dispersion estimate is typically lower. The use of the simplified form of the 2-D mass estimator, with velocity dispersion below the half-light radius, results in selecting parts of the velocity dispersion profile which are typically flat (aside from Draco), thus leading to lower velocity dispersion measurements and masses in the ‘light’ potential and higher in the ‘heavy’ potential. This trend is particularly prominent at dwarf apocentres in Figure 5, where the dwarfs are typically more spherical in 2-D and 3-D, thus minimizing the effects of any inaccuracies when estimating the half-light radius, or loss of equilibrium at pericentre.
We note that several sources in the literature report that rising line-of-sight velocity dispersion can be interpreted as the consequence of tides (Johnston et al., 1999; Read et al., 2006; Muñoz et al., 2008) due to inclusion of unbound stars in the tidal tails, however in our initial conditions this is simply a consequence of modelling a Plummer stellar distribution in an NFW halo121212See also Stoehr et al. 2002 for the velocity dispersion corresponding to the observed stellar density profiles embedded in cosmological NFW-like haloes and Peñarrubia et al. 2008 for the velocity dispersion of isotropic King profiles residing in NFW haloes, the Sculptor model of Iorio et al. 2019 (who also use a Plummer model for the stellar population) also produces a similar form of the line-of-sight velocity dispersion profile. and is particularly prominent in Fornax, Carina and Sculptor (see the Appendix). As a consequence of tides, the line-of-sight velocity dispersion profiles in our simulations tend to flatten and, in a heavy ‘potential’, drop towards the outskirts of the galaxy when only bound stellar particles are considered.
@@
This opens an interesting avenue to explore in constraining the mass of the Milky Way by jointly analysing the line-of-sight velocity dispersion profiles of its satellite dwarf galaxies. However, it would be of essence to obtain data beyond the half-light radii of the dwarf satellites, into their stellar halos (Deason et al., 2022; Jensen et al., 2024). We caution, however, that these conclusions also depend on the internal shape of the galaxies, their stellar light distribution (for a fixed model of dark matter), and full tidal history (since orbits of many dwarfs are affected by the presence of the LMC or may have been previously tidally stripped as part of group infall). More importantly, as we show in the Supplementary Materials, the presence of unbound stars in observational samples, which are difficult to remove through standard methods like 3 clipping, would make it hard to distinguish the case where the velocity dispersion rises due to weak effects of tides on the stellar profile from the one where the velocity dispersion rises because of the presence of unbound stars, indicative of stronger tidal effects.
@@ 3.1.4 Discussion
In a recent work, Errani et al. (2025) presented models of adiabatic and impulsive perturbations on stellar populations residing in dark matter haloes. The former case represents the effect on the stellar populations as the dwarf approaches a pericentre of an eccentric orbit. They found that, in cuspy haloes, an impulsive perturbation results in a series of damped oscillations in the half-light radius and the velocity dispersion, which further results in oscillations in the performance of a Wolf-like mass estimator that gradually decay. This is in agreement with the pattern we saw in Section 3.1.1, with an additional complication of minor changes in the axis ratio of the dwarfs, resulting in additional peaks in the performance of the estimator, particularly when the dwarf is viewed along the major axis. Errani et al. (2025) also find that in cored dark matter haloes the oscillations have a larger amplitude and are more long-lived, suggesting that it is worth exploring the performance of the estimator for individual dwarf galaxies in a similar setup to ours, but modelling the dark matter distribution with a cored profile.
Although we have simulated spherical systems, we found that tidal forces can cause deformations in the dwarfs’ shape. This effect, which is the most significant at pericentre, is clear in the dark matter but can also be seen in the stars, especially for eccentric dwarf galaxy orbits. The estimator recovers true within its errors in most cases and is typically accurate within 10 per cent, consistent with earlier findings of Errani et al. 2018 who applied the estimator to -body counterparts of Milky Way’s satellites derived from cosmological simulations. Nevertheless, our findings suggest the results of the estimator should be interpreted with care and to bear in mind the orbital properties and the current orbital phase of the subject galaxy. Additionally, it is unclear how the estimator would behave if the subject galaxy is aspherical and acted upon by tides. Our experiment with using dark matter as tracers demonstrates that for a population of tracers that extends to the tidal radius, the performance of the estimator tends to degrade, with mass typically underestimated by 15-20 per cent until the halo stops losing significant fractions of mass, becomes more spherical, and settles towards the final bound remnant state (Errani & Navarro, 2021).
@@ 3.2 Results with pyGravSphere
3.2.1 Effects of modelling assumptions
Fig. 8 show the results for Carina and Fornax, which have relatively low orbital eccentricity in both potentials, with orbits close to circular. For this reason, these galaxies have bound stars extending further out than the rest of our sample, with 8 kpc for Fornax and 2 kpc for Carina. While this is further out than the samples of data used by Read et al. (2019), in Supplementary Materials, we show that limiting the maximum projected radius of data produces consistent results. For Fornax in the ‘light’ and ‘heavy’ potential, the density is underestimated in the inner regions and overestimated near the half-light radius (black dashed line). This introduces a pattern that we will see in the remainder of our results. We note that such behaviour has been seen in the GravSphere method before when modelling Plummer-like stellar distributions inside cuspy haloes, particularly with a high number of stellar tracers (Read & Steger, 2017; Read et al., 2021). We have confirmed that this is not due to the permitted range of dark matter density log-slopes . We have tested this by modifying the priors of the broken power-law model to allow slopes as steep as . While this improved the fit to the outer profile in dwarfs that are particularly affected by tides, the fit to the inner profile did not improve, and the bias near the half-light radius remained. Truncating the sample of stellar particles to the inner 2 kpc in the case of Fornax did not significantly improve the fit in the ‘light’ potential in the outer regions of the halo, with the bias in the inner slope remaining. In the ‘heavy’ potential, however, this improved the recovery of the central density (see Supplementary Materials). We also found that this effect is unlikely due to the disequilibrium of the dwarfs, as this result persists whether the dwarf is modelled at pericentre, apocentre, or in isolation. We discuss the implications of this result in Section 3.3. In Appendix C we discuss in detail the origin of this bias, which is related to the inability of the broken power-law model used in pyGravSphere to describe the shape of the dark matter profile when the stellar population is significantly embedded within the dark matter halo (). This is due to the model’s parameterisation combined with the constraint on the “smoothness” of the broken power-law density profile.